find the center and radius for this circle
x^2-4x+y^2+2y=-1

Answer by stanbon(48535)

find the center and radius for this circle
x^2-4x+y^2+2y=-1
----------------
Complete the square on the x-terms and on the y-terms separately.
x^2-4x + 4 + y^2+2y + 1 = -1+4+1
(x-2)^2 + (y+1)^2 = 4
Center: (2,-1)
Radius: 2
===============
Cheers,
Stan H.